To calculate the angle of diffraction for red colour(λ=8000 angstrom)in the first order of spectrum for a grating with 14438 lines per inch.
1
Expert's answer
2019-11-27T10:16:39-0500

Condition for "m^{th}" interference maximum is "d sin \theta = m \lambda", where "d" is slit separation and "\lambda" is the wavelength.

We are given grating "N = 14438\frac{lines}{inch}", from where the slit separation is "d = \frac{1}{N} = 6.926 inch= 1.759 \cdot 10^{-6} m". The wavelength is "\lambda = 8000A = 800nm".

Hence, for the first maximum, "\sin \theta = \frac{\lambda}{d} \approx 0.455", which corresponds to "\theta \approx 27^{\circ}".